OpenStudy (anonymous):

When you write 1,000,000 as a power of 10: 1- the base is 1 and the exponent is 7. 2-the base is 1 and the exponent is 6. 3-the base is 10 and the exponent is 7. 4-the base is 10 and the exponent is 6.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

10^1=10 10^2=100 10^3=1000

OpenStudy (anonymous):

Can you see the pattern?

OpenStudy (anonymous):

So what is 10^7?

OpenStudy (anonymous):

10,000,000

OpenStudy (anonymous):

10 is the base and 7 is the power

OpenStudy (anonymous):

10^7=10,000,000

OpenStudy (anonymous):

thank you soo much

OpenStudy (anonymous):

and what is 10^6?

OpenStudy (anonymous):

10^6=1,000,000

OpenStudy (anonymous):

I just typed it into my calculator.

OpenStudy (anonymous):

You got your answer

geerky42 (geerky42):

another trick is to count zeros. \(\large 10^3 = 1,000\) note that 1,000 has three 0s? same thing for 10^7=10,000,000

OpenStudy (anonymous):

Forget the calculator. See the pattern.

geerky42 (geerky42):

10^10 = 10,000,000,000

OpenStudy (anonymous):

I see it but calculators don't lie.